# A new Linear Time Bi-level $\ell_{1,\infty}$ projection ; Application to the sparsification of auto-encoders neural networks

### v1

@@ 5.3.2 Experimental accuracy results on autoencoder neural neworks
Figure 10 shows the impact of the radius ($\eta$) on HIF2 dataset using by the bilevel projection.
From Tables 2 on synthetic dataset and Table 2 it can be seen that the best accuracy is obtained for $\eta=0.5$ for $\ell_{1,\infty}$ and for $\eta=1.$ for the bilevel $\ell_{1,\infty}$ projection, maximum accuracy of both method are similar. However, sparsity and computation time are better for bilevel $\ell_{1,\infty}$ than for regular $\ell_{1,\infty}$.
Synthetic 64 Baseline $\ell_{1,\infty}$ bi-level $\ell_{1,\infty}$
Best Radius 1 2.0
@@
Table 2: Synthetic dataset 64 features. SiLU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bi-level $\ell_{1,\infty}$.
Synthetic 16 Baseline $\ell_{1,\infty}$ bi-level $\ell_{1,\infty}$
Best Radius 0.5 1.0
[…]
Table 3: Synthetic dataset 16 features. SiLU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bi-level $\ell_{1,\infty}$.
Table 2 shows accuracy classification for 64 informative features. The baseline is an implementation that does not process any projection. Compared to the baseline, the SAE using the $\ell_{1,\infty}$ projection improves the accuracy by $10.3\%$.
Table 3 shows that accuracy results for 16 informative features of the bi-level $\ell_{1,\infty}$ and classical $\ell_{1,\infty}$ are similar. Again, compared to the baseline, the SAE using the bilevel $\ell_{1,\infty}$ projection improves the accuracy by $17.7\%$.
Real data HIF2 Baseline $\ell_{1,\infty}$ bi-level $\ell_{1,\infty}$
Best Radius 0.1 0.25
[…]
Table 4: HIF2 SiLU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bi-level $\ell_{1,\infty}$.
Again, on this real dataset, compared to the baseline the SAE using the bilevel $\ell_{1,\infty}$ projection improves the accuracy by $10\%$ similar to the results on the synthetic data-64.
6 Discussion
@@
Second, our bilevel method can be extended straight forward to tensor and convolutionnal neural networks. Thus, sparsity improves computational cost. Note that the $\ell_{1,1}$ projection was also used in order to sparsify convolutionnal neural networks [47] for green AI. This $\ell_{1,1}$ projection was also used for image compression using autoencoder [48] improving computational cost of new image coding method [37, 49] Our new bilevel $\ell_{1,\infty}$ projection will outperform all other projections in terms of computation cost and sparsity.
7 Conclusion
Although many projection algorithms were proposed for the projection of the $\ell_{1,\infty}$ norm, complexity of these algorithms remain an issue. The worst-case time complexity of these algorithms is $\mathcal{O}\big{(}n\times m\times\log(n\times m)\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. In order to cope with this complexity issue, we have proposed a new bi-level projection method. The main motivation of our work is the direct independent splitting made by the bi-level optimization, which takes into account the structured sparsity requirement. We showed that the theoretical computational cost of our new bi-level method is only $\mathcal{O}\big{(}n\times m\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. Experiments on synthetic and real data show that our bi-level method is 2.5 times faster than the actual fastest algorithm provided by Chu et. al.. Sparsity of our bi-level $\ell_{1,\infty}$ projection outperforms other bi-level projections $\ell_{1,1}$ and $\ell_{1,2}$.

### v2

@@ V-C2 Experimental accuracy results on autoencoder neural networks
Figure 8 shows the curve of the accuracy versus the radius ($\eta$) on HIF2 dataset. We can see that the curve is more well shaped using by the bilevel projection.
Real data HIF2 Baseline $\ell_{1,\infty}$ bilevel $\ell_{1,\infty}$
Best Radius - 0.1 0.25
@@
TABLE IV: HIF2 dataset SILU activation, Accuracy : comparison of $\ell_{1,\infty}$ and bilevel $\ell_{1,\infty}$.
Again, tableIV on this HIF2 real dataset, compared to the baseline the SAE using the bilevel $\ell_{1,\infty}$ projection improves the accuracy by $10\%$ similarly to the results on the synthetic data-64. Moreover, on this real dataset, our bilevel projection outperforms the usual $\ell_{1,\infty}$ projection by $1.2\%$.
Figure 9: Weights of the first layer of the Fully connected Neural network
Constraint optimization generally produces networks with random sparse connectivity, i.e. sparse weight matrices. They only reduce the memory cost but not the computational cost, since they result in networks with sparse weight matrices. Decreasing both memory and computational requirements can however be achieved by suppressing columns (features) instead of weights. Figure 9 shows that our bilevel $\ell_{1,\infty}$ suppress columns and thus efficiently reduces the computational cost.
VI Discussion
@@
Second, our bilevel method can be extended straight forward to tensor and convolutionnal neural networks. The $\ell_{1,\infty}$ projection has been successfully applied to the sparsification of autoencoders using convolutional neural networks with a memory reduction of 84 and a reduction in computational cost of 40 without visual image degradation [46] in the new image compression standard JPEG AI [47].
According to estimates, by 2040 Artificial Intelligence (AI) server farms may account for $14\textperthousand$ of all global carbon emissions in the world. Third, another application of our method can be the sparsification of attention matrices of a transformer architecture [48] used in AI software.
VII Conclusion
Although many projection algorithms were proposed for the projection of the $\ell_{1,\infty}$ norm, complexity of these algorithms remain an issue. The worst-case time complexity of these algorithms is $\mathcal{O}\big{(}n\times m\times\log(n\times m)\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$. In order to cope with this complexity issue, we have proposed a new bi-level projection method. The main motivation of our work is the direct independent splitting made by the bi-level optimization, which takes into account the structured sparsity requirement. We showed that the theoretical computational cost of our new bi-level method is only $\mathcal{O}\big{(}n\times m\big{)}$ for a matrix in $\mathbb{R}^{n\times m}$ and thus improves by $\mathcal{O}\big{(}\log(nm)\big{)}$ times classical algorithms. Note that this improvement can be huge for a large dataset. Experiments on synthetic and a real data show that our bi-level method is faster than the actual fastest algorithm. Sparsity of our bi-level $\ell_{1,\infty}$ projection outperforms other bi-level projections $\ell_{1,1}$ and $\ell_{1,2}$.